Grade 12th passGeneral Physics

Calculate the orbital velocity of satellite required moving around the earth where radius of earth is 6000000m and value of g is 10 m/s2

Feroz alam
7 Years agoGrade 12th pass

1 Answer

Rajat
7 Years ago
Orbital velocity=√(GM/R)=√gR=√10×6000000=
103×√60= 7.75×10m/s
I hope this is helpful to you.
With regards